# An Explicit Entire Function of Order One with All Zeros on a Line and Bounded in a Half-Plane

### v1

@@ 3. The Spectral Identity
$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$.
@@
Since $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term), the Euler–Maclaurin formula converts the zero sum to an integral: $\log|\Xi_{0}(s)|=\sum_{n}\log|1+w^{2}/(\gamma_{n}^{(0)})^{2}|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$. By Stirling’s formula, $\int\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du=\log|\Gamma_{\pi}(s)|+O(1)$ where $\Gamma_{\pi}(s)=\tfrac{1}{2}s(s-1)\pi^{-s/2}\Gamma(s/2)$ (the integral reproduces the Weierstrass product of $\Gamma$ at the Weyl density). Hence $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$. Note that $Z_{0}$ has a simple pole at $s=1$ from $1/[\tfrac{1}{2}s(s-1)]$; the bound is for $|t|\geq 2$. ∎
Remark 3.3.
Consequently, $\log|Z_{c}(\sigma+it)|=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(1)$ for $\sigma\geq 1$, $|t|\geq 2$.
4. The Fourier-Analytic Mechanism
The bounded/unbounded transition has a simple Fourier-analytic explanation. The spectral identity (Proposition 3.1) gives $\log|Z_{c}/Z_{0}|(\sigma+it)=-c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$. For $\sigma>1$: the sum $\sum\alpha_{k}^{-\sigma}$ converges (since $\alpha_{k}=p_{k}$ and $\sum p^{-\sigma}<\infty$), so the triangle inequality gives $|\log|Z_{c}/Z_{0}||\leq c\sum\alpha_{k}^{-\sigma}+O(c)$—boundedness. At $\sigma=1$: the sum $\sum\alpha_{k}^{-1}$ diverges (Mertens [11, Thm. 2.7]), and the Dirichlet series structure of $Z_{c}/Z_{0}$ (Remark 6.1) enables a Phragmén–Lindelöf argument (§5) to transfer the divergence into unboundedness of $Z_{c}$ on $\sigma=1$. The transition is thus driven entirely by the marginal convergence/divergence of the damped Fourier amplitudes $\sum\alpha_{k}^{-\sigma}$ at $\sigma=1$.
@@ 5. Proof of the Main Theorem
The function $\Xi_{c}$ is the Hadamard product (2) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via the full series $E$ as in (5). By Lemma 2.2, $\Xi_{c}$ is an entire function of order $1$.
@@
Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$ because the shifts $\delta_{n}=O(c\gamma_{n}^{(0)}/\log\gamma_{n}^{(0)})$ (Lemma 2.2) can change $N_{c}(T)$ by at most $O(\log T)$ relative to $N_{\mathrm{main}}(T)$.
Part (d) (Boundedness for $\sigma>1$). By Lemma 1.2, the linearization error satisfies $\sum\delta_{n}^{2}/(\gamma_{n}^{(0)})^{2}=O(c^{2})$. The spectral identity (Proposition 3.1) therefore holds for $\Xi_{c}$ with $O(c)$ error, giving
$\log|Z_{c}(\sigma+it)|\leq c\,\zeta_{\alpha}(\sigma)+O(1)<\infty$
for all $\sigma>1$, $|t|\geq 2$ (using $\log|Z_{0}|=O(1)$ from Remark 3.2).
Part (e) (Unboundedness on $\sigma=1$). The ratio $Z_{c}/Z_{0}$ is a Dirichlet series with generalized Euler product (Remark 6.1), absolutely convergent for $\sigma>1$, hence bounded on $\sigma=2$. We show it is unbounded on $\sigma=1$ by a Phragmén–Lindelöf argument (cf. [14, Thm. 8.4]).
@@
Step 1. $Z_{c}/Z_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$. For any $N$, by Dirichlet’s approximation theorem there exists $t\leq(2\pi/\varepsilon)^{N}$ with $|\omega_{k}t+\phi_{k}-\pi|<\varepsilon\pmod{2\pi}$ for all $k\leq N$. At $\sigma=1+1/\log N$:
$\operatorname{Re}\log\frac{Z_{c}}{Z_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{3}\log\log N-O(1),$
which grows without bound as $N\to\infty$ (since $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-2}\sum_{p\leq N}p^{-1}\sim e^{-2}\log\log N$ by Mertens).
@@
Step 2. Suppose $|Z_{c}/Z_{0}(1+it)|\leq M$ for all $t$. By Phragmén–Lindelöf in the strip $1\leq\sigma\leq 2$ (the function has at most polynomial growth on $\sigma=1$ and is bounded on $\sigma=2$), $Z_{c}/Z_{0}$ would be bounded throughout the strip. This contradicts Step 1.
@@
Therefore $Z_{c}/Z_{0}$ is unbounded on $\sigma=1$. Since $|Z_{0}(1+it)|=O(1)$ (Remark 3.2) and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$, it follows that $\sup_{t}|Z_{c}(1+it)|=+\infty$.
The transition is thus: for $\sigma>1$, $|Z_{c}|\leq\exp(c\,\zeta_{\alpha}(\sigma)+O(1))$ (unconditional by Lemma 1.2); on $\sigma=1$, $\sup_{t}|Z_{c}(1+it)|=+\infty$ (unconditional by the Dirichlet–PLF argument above). (Note: $Z_{c}$ has a pole at $s=1$ from the factor $1/[\tfrac{1}{2}s(s-1)]$; the large values are for $|t|\to\infty$.) ∎
6. The Ratio $Z_{c}/Z_{0}$ and Dirichlet Series
Remark 6.1 (Dirichlet series structure).
Since $\alpha_{k}=p_{k}$, the spectral identity becomes
$\log\frac{Z_{c}}{Z_{0}}(s)=-c\sum_{k=1}^{\infty}e^{-i\phi_{k}}\,p_{k}^{-s}+O(c),$
which is an ordinary Dirichlet series supported on primes. Exponentiating:
@@ 6. The Ratio $Z_{c}/Z_{0}$ and Dirichlet Series
(9) $\frac{Z_{c}}{Z_{0}}(s)=\prod_{p}\exp\!\bigl(-c\,e^{-i\phi_{k}}\,p^{-s}\bigr)=\sum_{n=1}^{\infty}a_{n}\,n^{-s},$
@@
with $|a_{p}|=c$,   $|a_{p^{m}}|=c^{m}/m!$,   and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$. This converges absolutely for $\sigma>1$. Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$ with the same $\Gamma/\pi$ prefactor, the ratio inherits a functional equation:
$\frac{Z_{c}}{Z_{0}}(s)=\frac{Z_{c}}{Z_{0}}(1-s).$
The ratio $Z_{c}/Z_{0}$ is meromorphic: it has zeros at $\{1/2+i\gamma_{n}^{(c)}\}$ (from $Z_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $Z_{0}$), both on $\sigma=1/2$. It is a Dirichlet series (with generalized Euler product) that is bounded for $\sigma>1$ and unbounded on $\sigma=1$, with all its zeros provably on the critical line.
The construction separates the Euler product (which lives on the ratio) from the entireness (which lives on $\Xi_{c}$). Note that this variant uses $\alpha_{k}=p_{k}$ (rational frequencies) and produces a single function $\Xi_{c}$ with the same properties as in Theorem 1.1; the ratio $Z_{c}/Z_{0}$ is then a Dirichlet series, but $Z_{0}$ is not.
@@ 7. Almost-Periodicity
The transition at $\sigma=1$ has a natural interpretation in the hierarchy of almost-periodic function spaces.
@@
    For $\sigma>1$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Bohr almost-periodic.
@@
    For $\sigma>\tfrac{1}{2}$: the function $t\mapsto\log|Z_{c}/Z_{0}|(\sigma\!+\!it)$ is Besicovitch ($B^{2}$) almost-periodic with
    $\|{\log|Z_{c}/Z_{0}|}\|_{B^{2}}^{2}=\tfrac{1}{2}\,c^{2}\sum_{k=1}^{\infty}\alpha_{k}^{-2\sigma}<\infty.$
@@
    The construction $E(T)$ is not $B^{2}$-AP: $\sum|a_{k}|^{2}=\frac{1}{\pi^{2}}\sum\alpha_{k}^{-1}=\infty$ (Mertens). The truncation $E_{K}\in B^{2}\text{-AP}$ with $\|E_{K}\|_{B^{2}}^{2}=\frac{1}{2\pi^{2}}\sum_{k=1}^{K}\alpha_{k}^{-1}$.
@@
(i) By Proposition 3.1, the function equals an absolutely convergent trigonometric series $c\sum\alpha_{k}^{-\sigma}\cos(\omega_{k}t+\phi_{k})+O(c)$ for $\sigma>1$. Since $\sum\alpha_{k}^{-\sigma}$ converges for $\sigma>1$ [11, Thm. 2.7], the series defines a Bohr-AP function.
(ii) The Fourier–Bohr coefficients at exponents $\pm\omega_{k}$ have squared amplitudes $(c\,\alpha_{k}^{-\sigma}/2)^{2}$. Since $\sum\alpha_{k}^{-2\sigma}$ converges for $2\sigma>1$, the Riesz–Fischer theorem for $B^{2}$-AP functions gives the result.
(iii) The amplitudes of $E(T)$ are $a_{k}=1/(\pi\alpha_{k}^{1/2})$. Since $\sum|a_{k}|^{2}=(1/\pi^{2})\sum\alpha_{k}^{-1}\to\infty$ (by Mertens’ theorem [11, Thm. 2.7]), $E$ does not lie in $B^{2}\text{-AP}$ (its mean square diverges). For the truncation $E_{K}$, the finite sum confirms $E_{K}\in B^{2}\text{-AP}$. ∎
The AP hierarchy for $Z_{c}$ mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the passage from convergent to divergent Fourier norm at $\sigma=1$ being the “Bohr-AP to $B^{2}$-AP” transition. The transitions occur at the same critical $\sigma$ values because $\alpha_{k}\sim p_{k}$; the difference is the source: arithmetic (Euler product) for $\zeta$, purely analytic (spectral identity) for $Z_{c}$.
@@ 8. Observations
We illustrate the construction numerically with $c=0.05$ and a finite truncation $K=80$. (The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ with $\theta=e^{\sqrt{2}}$, which ensures global monotonicity of $F_{K}$ (Remark 2.3); the qualitative behavior is identical to the $\alpha_{k}=p_{k}$ case.)

### v2

@@ 3.1. Linearization convergence
(11) $\sum_{n=1}^{\infty}\frac{E(\gamma_{n}^{(0)})^{2}}{n^{2}}<\infty.$
@@
Head (large sieve). $E_{K_{j}}$ is a trigonometric polynomial with $K_{j}$ terms and frequency range $\Lambda_{j}=\omega_{K_{j}}-\omega_{1}=O(j)$. The zeros $\gamma_{n}^{(0)}$ have minimum spacing $\Delta\geq c_{1}/j$ for $n\leq 2^{j+1}$. By the large sieve inequality for trigonometric polynomials at well-spaced points (cf. [13, Thm. 1], [14, §7]):
$\sum_{n\leq 2^{j+1}}\!|E_{K_{j}}(\gamma_{n}^{(0)})|^{2}\leq(\pi/\Delta+\Lambda_{j})\,\tfrac{1}{\pi^{2}}\sum_{k\leq K_{j}}p_{k}^{-1}\leq C_{1}\,j\log j,$
since $\sum_{k\leq K_{j}}p_{k}^{-1}\sim\log\log p_{K_{j}}\sim\log j$ (Mertens). Therefore $\text{HEAD}_{j}\leq C_{1}\,j\log j\,/\,4^{j}$, and $\sum_{j}j\log j/4^{j}<\infty$.
Tail (Abel summation). The tail $\sum_{k>K}(\omega_{k+1}-\omega_{k})/p_{k}^{1/2}$ is a left Riemann sum for the decreasing function $e^{-\omega/2}$. By Bertrand’s postulate (mesh width $\leq\log 2$): $\sum_{k>K}(\omega_{k+1}-\omega_{k})/p_{k}^{1/2}\leq\sqrt{2}\int_{\omega_{K}}^{\infty}e^{-\omega/2}\,d\omega=2\sqrt{2}\,p_{K}^{-1/2}$. Hence $|R_{K_{j}}(\gamma_{n}^{(0)})|\leq C_{2}\,2^{j(1-\varepsilon)/2}/j^{3/2}$ (with $\varepsilon=0.1$, using $\gamma_{n}\leq C_{3}\,2^{j}/j$ and $\sqrt{p_{K_{j}}}\geq C_{4}\,2^{(1+\varepsilon)j/2}\sqrt{j}$). Therefore $\text{TAIL}_{j}\leq C_{5}/(j^{3}\cdot 2^{\varepsilon j})$, and $\sum_{j}1/(j^{3}\cdot 2^{\varepsilon j})<\infty$.
Both contributions decay exponentially, so $\sum S_{j}<\infty$. ∎
3.2. The base function $Z_{0}$
Lemma 3.3.
$\log|Z_{0}(\sigma+it)|=O(1)$ uniformly for $\sigma\geq 1$ and $|t|\geq 2$. In particular, $|Z_{0}|$ is both bounded and bounded away from zero in this region.
Proof.
By construction, $\Xi_{0}$ has zero-counting function $N_{0}(T)=N_{\mathrm{main}}(T)$ (no error term). By Euler–Maclaurin, $\log|\Xi_{0}(s)|=\int_{\gamma_{1}^{(0)}}^{\infty}\log|1+w^{2}/u^{2}|\,N_{\mathrm{main}}^{\prime}(u)\,du+O(1)$. Stirling’s formula identifies this integral with $\log|\Gamma_{\pi}(s)|+O(1)$, giving $\log|Z_{0}|=\log|\Xi_{0}/\Gamma_{\pi}|=O(1)$. The second claim follows since $|\!\log|Z_{0}|\!|\leq C$ implies $e^{-C}\leq|Z_{0}|\leq e^{C}$. ∎
@@ 4. Proof of the Main Theorem
Parts (a)–(c) hold by construction; (d) follows from the spectral identity (Proposition 3.1) with linearization error controlled by Lemma 3.2; and (e) via a Phragmén–Lindelöf argument below.
The function $\Xi_{c}$ is the Hadamard product (6) with zeros $\gamma_{n}=\gamma_{n}^{(0)}+\delta_{n}$ defined via $E$ as in (5). By Lemma 2.1, $\Xi_{c}$ is an entire function of order $1$.
@@
Part (c): the counting function $N_{c}(T)=\#\{n:\gamma_{n}\leq T\}$ satisfies $|N_{c}(T)-N_{\mathrm{main}}(T)|=O(\log T)$. Indeed, zeros near height $T$ have index $n\sim T\log T/(2\pi)$ and perturbation $|\delta_{n}|=O(T^{1/2+\varepsilon}/\log T)$ by (2). In an interval of this length around $T$, $N_{\mathrm{main}}$ changes by $O(N_{\mathrm{main}}^{\prime}(T)\cdot T^{1/2+\varepsilon}/\log T)=O(T^{1/2+\varepsilon}/\log^{2}T)=o(\log T)$.
Part (d) (Boundedness for $\sigma>1$). The spectral identity (Proposition 3.1) gives
$\log|Z_{c}(\sigma+it)|=\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(s)+\log|Z_{0}(s)|\leq c\,P(\sigma)+O(1)$
for all $\sigma>1$, $|t|\geq 2$ (using $|\sum p_{k}^{-\sigma}\cos(\cdots)|\leq P(\sigma)$ and Lemma 3.3).
Part (e) (Unboundedness on $\sigma=1$). We show $\sup_{t}|\Xi_{c}/\Xi_{0}(1+it)|=+\infty$ by a Phragmén–Lindelöf argument (cf. [17, Thm. 8.4]).
@@
Step 1. $\Xi_{c}/\Xi_{0}$ is unbounded in $\{\sigma>1,\;|t|\geq 2\}$. For any $N$, by Dirichlet’s approximation theorem there exists $t\leq(2\pi/\varepsilon)^{N}$ with $|\omega_{k}t+\phi_{k}-\pi|<\varepsilon\pmod{2\pi}$ for all $k\leq N$. At $\sigma=1+1/\log N$:
$\operatorname{Re}\log\frac{\Xi_{c}}{\Xi_{0}}(\sigma\!+\!it)\geq c\sum_{k=1}^{N}p_{k}^{-\sigma}-O(c)\geq\tfrac{c}{e}\log\log N-O(1),$
which grows without bound as $N\to\infty$ (since $p_{k}^{-1/\log N}=e^{-\log p_{k}/\log N}\geq e^{-1}$ for $p_{k}\leq N$, so $\sum_{p\leq N}p^{-1-1/\log N}\geq e^{-1}\sum_{p\leq N}p^{-1}\sim e^{-1}\log\log N$).
@@
Step 2. Suppose $|\Xi_{c}/\Xi_{0}(1+it)|\leq M$ for all $t$. Since $\Xi_{c}$ and $\Xi_{0}$ are entire of order $1$, $|\Xi_{c}(\sigma+it)/\Xi_{0}(\sigma+it)|\leq\exp(C|t|^{1+\varepsilon})$ for any $\varepsilon>0$ (the ratio is analytic in $\sigma\geq 1$ since all zeros of $\Xi_{0}$ lie on $\sigma=1/2$). The Phragmén–Lindelöf principle for the strip $1\leq\sigma\leq 2$—bounded on $\sigma=1$ (by assumption) and on $\sigma=2$ (by part (d)), with at most exponential growth—gives boundedness throughout. This contradicts Step 1.
@@
Therefore $\Xi_{c}/\Xi_{0}$ is unbounded on $\sigma=1$, and since $Z_{c}/Z_{0}=\Xi_{c}/\Xi_{0}$ and $|Z_{0}(1+it)|$ is bounded away from zero for $|t|\geq 2$ (Lemma 3.3), $\sup_{t}|Z_{c}(1+it)|=+\infty$.
Part (f) (Approximate Euler product). This is Proposition 5.1 below. ∎
5. The Ratio $\Xi_{c}/\Xi_{0}$ and Dirichlet Series
Proposition 5.1.
For $\sigma>1$, the ratio $\Xi_{c}/\Xi_{0}$ admits the approximate Euler product representation
@@ 5. The Ratio $\Xi_{c}/\Xi_{0}$ and Dirichlet Series
(12) $\frac{\Xi_{c}}{\Xi_{0}}(s)=\kappa_{0}\cdot\prod_{k=1}^{\infty}\exp\!\bigl(-c\,e^{-ik\theta}\,p_{k}^{-s}\bigr)\cdot M(s),$
where $\kappa_{0}=e^{cC_{1}}$ is an explicit nonzero constant (with $C_{1}=C_{1}(\theta)$ the non-oscillatory constant from (10)), and $M(s)$ is bounded analytic with $M(s)$ bounded for $\sigma>1/2$, $|t|\geq 2$. The Euler product $\prod\exp(-ce^{-ik\theta}p_{k}^{-s})=\sum a_{n}\,n^{-s}$ has $|a_{p_{k}}|=c$, $|a_{p_{k}^{m}}|=c^{m}/m!$, and $a_{mn}=a_{m}a_{n}$ for $\gcd(m,n)=1$.
Proof.
The spectral identity (8) gives $\log(\Xi_{c}/\Xi_{0})(s)=-cD(s)+R(s)$ with $R$ bounded. Write R(s)=R(2)+(R(s)R(2))R(s)=R(2)+(R(s)-R(2)); the constant $\kappa_{0}=e^{R(2)}$ is a nonzero number and $M(s)=e^{R(s)-R(2)}$ is bounded analytic. Then $\Xi_{c}/\Xi_{0}=\kappa_{0}\cdot\exp(-cD(s))\cdot M(s)$. Expanding the exponential of the prime-supported Dirichlet series gives the Euler product and multiplicative coefficients. ∎
@@
Since both $\Xi_{c}$ and $\Xi_{0}$ satisfy $\Xi(s)=\Xi(1\!-\!s)$, the ratio inherits the functional equation $\Xi_{c}(s)/\Xi_{0}(s)=\Xi_{c}(1-s)/\Xi_{0}(1-s)$. The ratio is meromorphic: it has zeros at $\{1/2+i\gamma_{n}\}$ (from $\Xi_{c}$) and poles at $\{1/2+i\gamma_{n}^{(0)}\}$ (from $\Xi_{0}$), all on $\sigma=1/2$.
@@ 6. Almost-Periodicity
The Dirichlet series $D(s)$ itself has a clean almost-periodicity hierarchy.
@@
    For $\sigma>1$: $t\mapsto D(\sigma+it)$ is Bohr almost-periodic (uniformly AP), since $D$ is an absolutely convergent Dirichlet series.
@@
    For $\sigma>\tfrac{1}{2}$: $t\mapsto D(\sigma+it)$ is Besicovitch ($B^{2}$) almost-periodic with $\|D(\sigma+i\,\cdot\,)\|_{B^{2}}^{2}=\sum_{k=1}^{\infty}p_{k}^{-2\sigma}<\infty$.
@@
    The restriction $E(T)=\pi^{-1}\operatorname{Im}\,D(\tfrac{1}{2}+iT)$ is not $B^{2}$-AP: $\sum p_{k}^{-1}=\infty$ (Mertens).
@@
(i) Absolute convergence of $\sum p_{k}^{-\sigma}$ for $\sigma>1$ gives uniform convergence of the trigonometric series, hence Bohr-AP. (ii) The Fourier–Bohr coefficients of $D(\sigma+it)$ at exponents $\omega_{k}=\log p_{k}$ have squared moduli $p_{k}^{-2\sigma}$. Since $\sum p_{k}^{-2\sigma}$ converges for $\sigma>1/2$, the Riesz–Fischer theorem gives the result. (iii) Since $\sum p_{k}^{-1}=\infty$ (Mertens), the Fourier energy diverges at $\sigma=1/2$. ∎
The spectral identity transfers these properties to $\log(\Xi_{c}/\Xi_{0})$ up to the bounded error $R(s)$. The Fourier–Bohr coefficients of $\log(\Xi_{c}/\Xi_{0})$ at frequencies $\omega_{k}$ are $c\,e^{-ik\theta}\,p_{k}^{-\sigma}$ to leading order; since $R$ is bounded but need not itself be almost-periodic, the AP properties of $\log(\Xi_{c}/\Xi_{0})$ hold to leading order.
The AP hierarchy mirrors that of $\zeta$: Bohr-AP for $\sigma>1$ (Bohr [1]), $B^{2}$-AP for $\sigma>1/2$ (Bohr–Jessen [3]), with the convergent-to-divergent Fourier norm transition at $\sigma=1$.
@@ 7. Observations
We illustrate the construction numerically with $c=0.05$ and a finite truncation $K=80$. The computations use the shifted variant $\alpha_{k}=p_{k}+\theta$ in place of $p_{k}$, which ensures global monotonicity of $F_{K}(T)=N_{\mathrm{main}}(T)+cE_{K}(T)$ (see Remark 7.1). Since $\alpha_{k}\sim p_{k}$ as $k\to\infty$, the convergence/divergence properties at $\sigma=1$ are unchanged, and the spectral identity has the same structure with $\alpha_{k}$ replacing $p_{k}$.
Remark 7.1.
For the finite truncation $E_{K}$, global monotonicity of $F_{K}$ holds when $c<N_{\mathrm{main}}^{\prime}(T_{0})/(\sum_{k=1}^{K}\omega_{k}\alpha_{k}^{-1/2})$, which is satisfied for the parameters used. For the full infinite series, occasional crossings ($\gamma_{n+1}<\gamma_{n}$) may occur but do not affect the Hadamard product or the spectral identity (which depend on the linearization at $\gamma_{n}^{(0)}$, not on the ordering of $\gamma_{n}$).